marcus and beth leave college traveling in opposite directions on a straight road beth drives 13 mi/h faster the marcus after 4 hours they are 452 miles apart find their rates
Answer by ikleyn(53765)

.

Let r be the Markus rate;  then the Beth rate is (r+13) mph.


The total separating distance is  4r + 4*(r+13) miles,

and it is equal to 452 miles.  So your base equation is

    4r + 4*(r+13) = 452  miles.


Simplify and find r

    4r + 4r + 52 = 6730

    8r           = 452 - 52 = 400

     r                      = 400/8 = 50 miles per hour.


ANSWER.  The Markus rate is  50 mph.

         The Beth rate is  50 + 13 = 63 mph.


CHECK.  4*50 + 4*63 = 452 miles separating distance in 4 hours.   ! Correct !

Solved.

-----------------

For simple Travel & Distance problems,  see introductory lessons
    - Travel and Distance problems
    - Travel and Distance problems for two bodies moving in opposite directions
    - Travel and Distance problems for two bodies moving in the same direction (catching up)
in this site.

They are written specially for you.

You will find the solutions of many similar problems there.

Read them and learn once and for all from these lessons on how to solve simple Travel and Distance problems.

Become an expert in this area.